OpenStudy (anonymous):

The table below shows the values of y for each value of x: x y -1 0 -1 -1 -2 2 -2 3 Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. Part B: The function f(h) shown below represents the number of tennis matches Diana won each month when she practiced tennis for h hours every day: If f(h) = 3h + 1 Calculate and interpret the meaning of f(3). Part C: Write an ordered pair to represent the input and output of the function in Part B when Diana practices tennis for b hours every

OpenStudy (anonymous):

@Compassionate @ganeshie8 @RBauer4 @TuringTest @ankitshaw @texaschic101

OpenStudy (anonymous):

Can you try to fix your table? I can't read it. I can help you with the other parts. For part b they are asking you to plug in 3 for h and solve it. f(3)=3(3)+1 do you understand ?

OpenStudy (anonymous):

for x -1 -1 -2 -2

OpenStudy (anonymous):

for y 0 -1 2 3

OpenStudy (anonymous):

Ok for the first one it is not a function because there are 2 values of y for 1 x.

OpenStudy (anonymous):

I would input -1 and get 2 values 0 and -1

OpenStudy (anonymous):

And did you calculate the value for f(3)?

OpenStudy (anonymous):

i got that already, this is what i have so far. All ordered pairs must have an x value. X is not a function.

OpenStudy (anonymous):

Ok yes all ordered pairs need to have an x value but it order for it to be a function there must be only one y value for each x value

OpenStudy (anonymous):

And since in yours the value -1 has both 0 and -1 as y values it is not a function

OpenStudy (anonymous):

f(h) = 3h + 1 have not gotten this

OpenStudy (anonymous):

Ok for that one you plug 3 in for h. f(3)=3(3)+1, what is 3*3?

OpenStudy (anonymous):

f(h) = 3h+1, so f(3) = 3(3)+1 = 10 she wins 10 matches if she practices 3 hours a day.

OpenStudy (anonymous):

Yes that is correct!

OpenStudy (anonymous):

Now to write and ordered pair for that, you had your input has 3 and the output was 10, x is the input and y is the output so it is (3,10)

OpenStudy (anonymous):

i dont understand the last part

OpenStudy (anonymous):

ohhh okay thank you!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Do you have time for 2 more?

OpenStudy (anonymous):

Sure, mind if you post in another question ?